OpenStudy (anonymous):

Find the length of the hypotenuse of a 45*,45*,90* triangle with a leg of 8.

OpenStudy (anonymous):

The choices are: a.\[4\sqrt{2}\] b.16 c.8 d\[8\sqrt{2}\]

OpenStudy (anonymous):

And this "*" Is used as the degree's sign

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@jordanloveangel @potatoprincess22

OpenStudy (studygurl14):

Hint: If the leg of a 45-45-90 triangle is 1...

OpenStudy (anonymous):

1.14

OpenStudy (studygurl14):

um...no

OpenStudy (studygurl14):

If the leg of a 45-45-90 triangle is 1, then the hypotenuse is \(\sqrt{2}\) If the leg of a 45-45-90 triangle is 8, then what is the hypotenuse?

OpenStudy (anonymous):

Would it be the \[\sqrt{16}\]?

OpenStudy (studygurl14):

No...It wouldn't

OpenStudy (studygurl14):

Set up this proportion: \(\large \frac{1}{\sqrt{2}}=\frac{8}{x}\) Solve for x

OpenStudy (studygurl14):

(cross multiply)

OpenStudy (wolf1728):

Here's a web page that will explain about 45 45 90 triangles http://www.1728.org/trig2.htm Anyway in a 45 45 90 triangle, the hypotenuse equals the leg times square root of 2 (about 1.414)

OpenStudy (wolf1728):

So if the leg = 8 then the hypotenuse equals?

OpenStudy (wolf1728):

Are you still there babygurl47672?

OpenStudy (studygurl14):

@babygurl47672

OpenStudy (anonymous):

Yeah I am sorry I'm so busy with all my work rn

OpenStudy (wolf1728):

So, hey ... how about those Boston Bruins? :-)

OpenStudy (wolf1728):

Anyway, the hypotenuse = leg * square root of 2 So the hypotenuse = 8 * 1.414 = 11.31

OpenStudy (anonymous):

the exact answer would be D, 8(square root of)2 because (square root of)2 is an irrational number, and therefore in order to get the exact solution you need to keep it as (square root of)2

OpenStudy (anonymous):

@babygurl47672 do you still need help?

OpenStudy (anonymous):

No I'm good thanks